You are a soybean farmer in Somewheregreat, Texas. You know that the own price elasticity of supply is 0.25 and own price elasticity of demand is -0.50. Last year the U.S. had a total demand of 4.3 billion bushels and produced 4.4 billion bushels of soybeans. Due to tariffs imposed by trading partners, the price of soybeans has dropped 20%. Find the new quantity demanded and the new quantity supplied for soybeans. Are soybean farmers better off? Show your work.
Answer #1

Elasticity of demand = -0.50

Elasticity of supply = 0.25

Quantity demanded = 4.3 billion

Quantity supplied = 4.4 billion

Price decreases by 20%

New quantity demanded: -0.50 = D/20, D = 20*0.5 = 10%

New quantity demanded decreases by 10% to 4.3 billion - 430000 = 3870000

New quantity supplied: 0.25 = S/10%, S= 10*0.25 = 2.5%

New quantity supplied increases by 2.5% to 4.4 billion+110000 = 4510000